# Utilizing smartphone sensors for accurate solar irradiance measurement and educational purposes

### v2

@@ 3 Materials and methods
The objective of this work is to demonstrate the usefulness of smartphones as an experimental measurement tool for direct normal solar irradiance. This requires mounting a tube around the smartphone light sensor, pointing it directly at the Sun, and then calibrating its measurement. After calibration, the smartphone can be used to determine parameters of the state of the atmosphere, such as direct transmittance, , or Linke’s turbidity, .
In this study, the calibration of the equipment is conducted using two distinct approaches: (i) by comparing it to high-quality pyrheliometer data obtained from professional measurements, and (ii) by comparing it to estimates from sophisticated publicly available clear-sky models. Both calibration methods require clear-sky conditions to ensure consistent measurements and to mitigate any discrepancies associated with cloud movement.
@@ 3.1 Experimental measurements
The precise measurement of DNI presents some difficulties. To carry it out, a pyrheliometer is used, an instrument that consists of an array of thermocouples (pyranometer) attached to a collimator tube, and a precision solar tracking mechanism. If the equipment is aligned with a precision of less than °, the pyrheliometer is capable of measuring the DNI with an uncertainty about 1. The measurements are carried out in broadband, that is, the irradiance corresponding to wavelengths between 200-4000 nm (which includes the entire solar spectrum) is integrated into a single value. Figure 2 shows the experimental setup of this work consisting of a Kipp & Zonen [22] CHP1 pyrheliometer (blue oval) and a Samsung S5 smartphone (yellow oval) assembled in a Solys2 precision solar tracker. The assembly of the smartphone is shown in Fig. 3, in particular, its location perpendicular to the axis of the black bars and the assembly of the small hand-made collimator tube for the light sensor.
@@
The measurements were made at the Solar Energy Laboratory (LES) of the University of the Republic (Udelar). The experimental site of this laboratory is located in Salto department, in northwestern Uruguay, with geographic coordinates of (latitude) and (longitude), corresponding to the UTC-3 time zone. The signal generated by the CHP1 pyrheliometer (in mV) is recorded by a Fisher Scientific DataTaker DT85 data logger and is converted to irradiance (in W/m) through the equipment constant. This measurement is the reference DNI measurement of the LES lab, and it is recorded continuously on a minute scale as an average of instantaneous measurements every 15 s. For the smartphone measurements, the ambient light sensor embedded in the Samsung S5 and the Phyphox [23] application are used. A diffuser is placed above the sensor, in this case, tracing paper printed in black, which prevents saturation of the recorded signal. A cylindrical tube painted black is also placed, which acts as a collimator for a large part of the diffuse irradiance, emulating the professional pyrheliometer collimator (see Fig. 3). Lighting measurements with the light sensor are recorded by the Phyphox app on a minute scale. It can also be on other time scales, for example 5-minute. Once the smartphone is placed in the solar tracker, the collection begins keeping the device measuring during the several hours. To protect the phone screen while the measurements are recorded throughout the day, a double sheet of white paper (A4) placed in front of the smartphone screen, acting as a radiation blocker to prevent the device from overheating as shown in Fig. 3.
3.2 Practical considerations for the measurements using smartphones
@@
The selection of an appropriate diffuser is crucial in achieving accurate measurements with a smartphone. The smartphone’s analog-digital converter incorporates internal electronics that adjust its gain based on the illuminance detected by the sensor. Consequently, if the solar radiation measurements are low (below klx, in this case), the equipment will automatically change its scale without notifying the user. Each scale is associated with a specific sensor saturation value, and this scaling behavior can result in erroneous measurements for significant periods when the measurement is in close proximity to the saturation value. Furthermore, for the specific smartphone used in this study, the upper limit of measurable illuminance in the higher range scale is klx, which represents the saturation threshold. As a point of reference, Michael et al. [24] obtained a conversion constant of lx / W m, indicating that measuring 1000 W/m would not be feasible with our smartphone (approximately klx) without the inclusion of a diffuser. Hence, it is essential to regulate the attenuation of illuminance before it reaches the sensor, for two primary reasons: (i) to ensure that values can be accurately recorded without saturating the sensor, and (ii) to maintain consistent measurements within a specific range of scales at all times. This implies that the introduced attenuation by the diffuser must strike a balance, neither being too minimal nor too excessive, but rather falling within an intermediate range.
@@
To illustrate the impact of different diffusers, we present the results obtained from two clear-sky days using two distinct types of diffusers, as depicted in Fig. 4. The diffusers employed were: (a) ordinary white paper with a surface mass density of 120 g/m, and (b) black printed tracing paper. The graph indicates the illuminance measurements captured by the smartphone (indicated in black) with each diffuser, alongside the reference direct irradiance measurements obtained from the pyrheliometer (displayed in blue), and the clear-sky satellite estimates (depicted in red). The behavior of the measurement obtained using diffuser (a) is illustrated in the left panel of Fig. 4, where the various scale changes occurring at low illuminance levels between and klx are clearly observed, along with the corresponding saturation points on each scale. A similar behavior at low illuminance can be observed for diffuser (b) in the right panel of Fig. 4, but only for values below klx, with notable prominence during sunset. For measurements within the range of values exceeding klx, the equipment does not undergo scale changes, resulting in continuous and seemingly anomaly-free measurements facilitated by diffuser (b). It is also evident from the graph that the ordinary white paper diffuser attenuates the signal to a greater extent compared to the tracing paper diffuser (as depicted on the right-hand side -axis of both plots). This difference can be attributed to the higher reflectivity of ordinary white paper, particularly within the visible region of the solar spectrum. Therefore, based on our findings, we recommend the use of diffuser (b) in this study. Custom selections may be done for other smartphones, however, this point requires special attention.
3.3 High precision clear-sky estimates
@@
As an alternative calibration method for places where a professional DNI measurement is not available, it is possible to use as reference accurate clear-sky estimates that use information from weather satellites and physically-consistent atmospheric models. This change in the reference implies a slight increase in the uncertainty in the determination of the calibration constants, since the DNI satellite estimate presents greater uncertainty than a ground reference measurement. There are sophisticated clear-sky models that integrate estimates of different atmospheric variables, either by satellite or by atmospheric reanalysis models, which can be considered as reference [13, 25] if they have been validated by terrestrial measurements with good concordance in various parts of the world.
@@
One interesting choice is the CAMS [26] (Copernicus Atmosphere Monitoring Service) platform which provides free clear-sky estimates using one of these reference models for the entire globe, the McClear model [25]. This model is based on sophisticated radiative transfer calculations from the libRadTran [27] library and its operational version takes the form of a multiple input table based on real-time information on the state of the atmosphere. In particular, this model uses information on aerosols, precipitable water column, and ozone obtained from the CAMS reanalysis database and Earth albedo estimates obtained by the MODIS low-orbit satellite. The CAMS reanalysis in turn assimilates weather satellite’s information to provide its modeled data. Its platform enables access to 1-minute (and other time scales) solar irradiance estimates (global, direct and diffuse) from this high-precision model by simply entering the latitude and longitude of the site of interest. The file header contains information on each of the solar magnitudes provided. For example, the dimensions of radiation are Wh m (irradiation, energy in the time interval per unit area), which must be converted to W m (average power per unit area) by the corresponding conversion depending on the time scale.
4 Calibration
@@
will be used, where is the illuminance measured by the smartphone expressed in klx, are the DNI data expressed in W m, and and are two conversion constants to adjust. The calibration is performed with the two reference DNI data sets considered, the professional measurement of the pyrheliometer and the estimates of the McClear model. Minute measurements and estimates of November 24, 2021 were used at the LES experimental site, where clear-sky conditions were maintained throughout the day. Observing the plot of panel b in Fig. 4, for the adjustment we used only the data that meet klx, so that the smartphone sensor was always on the same scale of measurement. This value was chosen conservatively, in order to ensure measurements at intermediate values on the scale. The calibration constants for both cases are presented in Table 1 and the experimental fit in Fig. 5.
@@
It is interesting to compare the calibration curves. Satellite estimates exhibit deviations from the terrestrial measurements, so the calibration curve based on these data will present more uncertainty than the one obtained by comparison with ground measurements. As observed in Fig. 4, the McClear estimates for that day overestimate the direct irradiance. This leads the calibration curve obtained with this data set results above the calibration curve obtained with terrestrial measurements as observed in Fig. 5. Comparison of the McClear calibration curve with the measured DNI data reports a mean deviation of +2.1% (overestimation) and a mean square deviation of 2.6%. This uncertainty is above the measurement uncertainty of the reference instrument (1%), so it is distinguishable, but at the same time it is below the typical uncertainty of clear-sky satellite models (3-6%) [28]. This demonstrates that it is possible to perform the calibration based on satellite data of solar irradiance with a low uncertainty, enabling its use in the absence of reference measurements. Figure 6 shows the measurements obtained with the smartphone using both calibrations. As can be seen, in both cases a very good DNI measurement is achieved and completely acceptable for an instructional laboratory for a wide range of the day. The only downside is that the smartphone does not achieve a good measurement in the first and last minutes of the day where the illuminance is very low. Outside the diurnal range, i.e. when the sensor measurement is zero, the DNI measurement is affected by the non-zero offset of Eq. 11.
@@
These experimental results reveal that the smartphone is an outstanding tool for measuring solar irradiance through its illuminance sensor, applicable in low-cost university physics laboratories. Furthermore, the measurement capacity achieved with the equipment is really good, even in comparison to commercial sensors. Two relevant questions regarding the measurement capacity of the smartphone arises here. The first is to evaluate the typical uncertainty of a smartphone sensor used to measure DNI with both calibration methods. To answer this question would require data acquisition for several consecutive days (2-3 weeks) similar to professional calibrations following international standards [21]. The second is to evaluate the stability of the calibration curve over time, that is, how robust is the sensor to gradual degradation. This would require a professional calibration of the sensor every 3 months for about a year. With this set of tests it is possible to technically evaluate the capacity limits of the smartphone sensor for the solar irradiance measurement. In fact, moderate-cost professional irradiance sensors are recommended to be calibrated once a year, and a similar recommendation can arise for the smartphone sensor. We recommend here to perform the smartphone calibration each time before its utilization (i.e. the first day of measurements). Similar studies can be carried out for the measurement of global irradiance in the horizontal plane, not only direct irradiance, which surely requires evaluating the non-planar angular response of the smartphone sensor.
5 Instructional laboratory
The preceding section outlined the calibration procedure for transforming a smartphone’s illuminance sensor into a reliable solar irradiance meter, utilizing both high-quality ground measurements and clear-sky satellite estimates. The latter option offers broader accessibility, making it applicable in diverse socio-educational contexts. Once the sensor is appropriately calibrated, it becomes a valuable tool for conducting various solar resource measurements, making it particularly useful in university laboratory settings. The calibrated smartphone can be employed to address numerous aspects related to solar radiation measurement. For instance, it enables the examination of solar resource variations across different geographical locations, under varying cloud conditions, and at different time scales (e.g., throughout a single day, consecutive days, or different seasons). Additionally, by performing measurements over the course of an entire day and integrating the obtained irradiance values, it becomes possible to estimate the daily irradiation. This opens up opportunities for conducting comprehensive analyses of solar radiation characteristics, for example, through a year.
In this section, our focus shifts towards presenting an instructional laboratory that focus into the physics of solar radiation and its interaction with the Earth’s atmosphere. This laboratory is designed around the measurement of a full day using the calibrated smartphone, enabling students to gain a deeper understanding of solar radiation attenuation and its intricate relationship with atmospheric interactions.
The first step consists of calculating the solar trajectory in the celestial sphere for the day in which the measurements will be made in order to know the position to which the smartphone should be tilted. The location of the Sun is determined with the angles of altitude, , and azimuth, , calculated by
and
As (latitude) is a fixed value and (solar declination) can be considered constant within a day, the intraday variation of these parameters is given by the solar zenith angle, which can be calculated using Eq. 2 based on a time vector specific to the day under consideration. The intraday varying magnitude is intended to correct a common sign artifact of the inverse function, and might not be needed depending on the computing framework. The graph of these two quantities allows to elaborate a solar diagram, as seen in Fig. 7, which illustrates the trajectory of the Sun in the sky for a given day. From this graph it is possible to plan the time to take measurements, for example, by identifying the local time of solar noon (maximum solar height). Additionally, the solar diagram aids in understanding the changing patterns of solar radiation throughout the day.
To conduct the experiment, additional variables need to be obtained, including the orbital factor, the extraterrestrial irradiance corrected by the orbital factor, the air mass (as defined in Eq. 7), and the global optical depth of the Rayleigh extinction (as defined in Eq. 9) for the specific measurement day. While the first two variables exhibit negligible variations within a single day, the latter two vary geometrically throughout the day. To capture these intraday variations, a time vector with minute-scale resolution or 10-minute scale is necessary.
In the subsequent analysis, we will utilize the calibrated measurements obtained from the smartphone over a full day with a minute resolution using the experimental setup illustrated in Fig. 2. In situations where a solar tracking system is not available, the measurements can be performed manually by aligning the smartphone (with the home-made collimator tube attached) towards the Sun and recording measurements at a chosen time interval (such as 10-minute or 15-minute intervals) for a time lapse of 1 to 2 hours centered around solar noon. In this experiment we focused on two activities:
1)
Determine the average Linke’s turbidity for the day of measurements, . The instantaneous value of this quantity (in this case, minute) is obtained from Eq. 14. To derive a single daily value of Linke’s turbidity for the entire day or subsequent days, the minute values are averaged around solar noon. This averaging process takes into account the stability of direct radiation during this time period, where its value remains relatively constant, as illustrated in Fig. 4. Therefore, the period around solar noon provides the most favorable conditions for making this determination, ensuring accurate and consistent measurements of turbidity throughout the day.
2)
Estimate the DNI for that day with the ESRA model and the estimated local , and compare it to the smartphone ground measurements. This comparison can include the satellite estimates against the device measurement for a comprehensive evaluation of accuracy and reliability.
In the next subsection we conduct a thorough assessment of accuracy and reliability by comparing the smartphone ground measurements with the ESRA model estimation and satellite estimates.
5.1 Determination of the average Linke’s turbidity
The determination of the Linke’s turbidity is based on the smartphone data using the linear relationship given by Eq. 11 averaged around solar noon. In this case, the most reliable linear regression corresponds to the measurements obtained with the reference pyrheliometer. Figure 8 displays the instantaneous values of from 10:00 to 15:00 local time. The samples highlighted in red are within hours of solar noon for that day (at 12:38). Solar noon’s time can be retrieved from several websites making these kinds of calculations, or by looking the time in which is maximum (or ). These samples are used to determine the average value of Linke’s turbidity, , with a standard deviation of 0.03, which falls within the expected range for a day with high atmospheric clarity.
5.2 ESRA model with experimental TL
After determining the average Linke’s turbidity it is possible to estimate the direct irradiance for that day with the ESRA model, according to
Figure 9 presents the ground-based DNI measurements, McClear model satellite estimates, and the ESRA model results for the specific day. It is evident that the simplified ESRA model slightly underestimates the clear-sky irradiance for this particular day. On the other hand, the McClear model, which is more complex and incorporates additional atmospheric input information, provides a better fit to the ground measurements, albeit with a slight overestimation. The ESRA model, despite its simplicity, effectively captures the overall behavior of clear-sky irradiance by considering, with only one parameter, various scattering and absorption effects related to solar radiation interaction with different clear-sky atmosphere components (e.g., air, water vapor, aerosols, stratospheric ozone). Considering its performance and ease of implementation, the ESRA model is recommended for inclusion in instructional laboratories about the solar resource. Furthermore, the Linke’s turbidity value can be fairly utilized for subsequent days to assess the behavior of solar irradiance (measured with the calibrated smartphone) in comparison to a clear-sky model adjusted by the students themselves. This allows for the observation and comparison of solar irradiance patterns on different clear-sky days or with varying levels of cloudiness (in cloudy days, for instance, the modeled clear-sky irradiance should be an upper limit of possible irradiance measurements). Additionally, it is possible to quantify the performance of models against ground measurements using the mean deviation and root mean square deviation. In particular, the CAMS website also provides all-sky satellite estimates that can be assessed by students with the calibrated smartphone.
It is common practice to calculate performance indicators by excluding samples with low solar elevations [29, 30] (such as the first and last samples of the day), typically exceeding 5-10 °, as depicted in Figure 9. Alternatively, this threshold could be set with another criterion such as klx. For instructional purposes, it is recommended to filter samples based on solar elevation () and with an exigent criterion, ensuring measurements quality. Table 2 presents the percentage performance metrics using the guideline. The calculated quantitative indicators for the study day confirm the underestimation of the ESRA model and the slight overestimation of the McClear model, both falling within the % range. As expected, the relative root mean square deviation (rRMSD) of the ESRA model is higher, but both models exhibit rRMSD values below 5%, which aligns with the expected performance range for clear-sky models.
| Model | rMBD (%) | rRMSD (%) |
| ESRA with averaged Linke’s turbidity. | % | % |
| McClear obtained with CAMS | % | % |

### v3

@@ 3 Materials and methods
The objective of this work is to demonstrate the usefulness of smartphones as an experimental measurement tool for direct normal solar irradiance. This requires mounting a tube around the smartphone light sensor, pointing it directly at the Sun, and then calibrating its measurement. In this study, the calibration of the equipment is conducted using two distinct approaches: (i) by comparing it to high-quality pyrheliometer data obtained from professional measurements, and (ii) by comparing it to estimates from sophisticated publicly available clear-sky models. Both calibration methods require clear-sky conditions to ensure consistent measurements and to mitigate any discrepancies associated with cloud movement.
@@ 3.1 Experimental measurements
The measurements were made at the Solar Energy Laboratory (LES) of the University of the Republic (Udelar). The experimental site of this laboratory is located in Salto department, in northwestern Uruguay, with geographic coordinates of (latitude) and (longitude), corresponding to the UTC-3 time zone.
@@
In this experiment, the ambient light sensor of smartphone Samsung S5 were used thanks to the freely available Phyphox app University (2020). A simple diffuser is placed above the sensor, in this case, tracing paper printed in black, which prevents saturation of the recorded signal. A cylindrical tube painted black is also placed, which acts as a collimator for a large part of the diffuse irradiance, emulating the professional pyrheliometer collimator (see Fig. 3). We emphasize that the selection of the diffuser offers no difficulties since its only function is to attenuate the radiation to avoid saturation and therefore does not affect the calibration procedure. Lighting measurements with the light sensor are recorded by the Phyphox app on a minute scale. Once the smartphone is placed in the solar tracker, the collection begins keeping the device measuring during a interval from several minutes to hours. To protect the phone screen while the measurements are recorded throughout the day, a double sheet of white paper (A4) placed in front of the smartphone screen, acting as a radiation blocker to prevent the device from overheating as shown in Fig. 3.
@@
The selection of an appropriate diffuser is important in achieving accurate measurements with a smartphone. The smartphone’s analog-digital converter incorporates internal electronics that adjust its gain based on the illuminance detected by the sensor. Consequently, if the solar radiation measurements are low (below klx, in this case), the equipment will automatically change its scale without notifying the user. Each scale is associated with a specific sensor saturation value, and this scaling behavior can result in erroneous measurements for significant periods when the measurement is in close proximity to the saturation value. Furthermore, for the specific smartphone used in this study, the upper limit of measurable illuminance in the higher range scale is klx, which represents the saturation threshold. As a point of reference, Michael et al. Michael et al. (2020) obtained a conversion constant of lx / W m-2, indicating that measuring 1000 W/m2 would not be feasible with our smartphone (approximately klx) without the inclusion of a diffuser. Hence, it is essential to regulate the attenuation of illuminance before it reaches the sensor, for two primary reasons: (i) to ensure that values can be accurately recorded without saturating the sensor, and (ii) to maintain consistent measurements within a specific range of scales at all times. This implies that the introduced attenuation by the diffuser must strike a balance, neither being too minimal nor too excessive, but rather falling within an intermediate range.
@@
To illustrate the impact of different diffusers, we present the results obtained from two clear-sky days using two distinct types of diffusers, as depicted in Fig. 4. The diffusers employed were: (a) ordinary white paper with a surface mass density of 120 g/m2, and (b) black printed tracing paper. The graph indicates the illuminance measurements captured by the smartphone (indicated in black) with each diffuser, alongside the reference direct irradiance measurements obtained from the pyrheliometer (displayed in blue), and the clear-sky satellite estimates (depicted in red). The behavior of the measurement obtained using diffuser (a) is illustrated in the left panel of Fig. 4, where the various scale changes occurring at low illuminance levels between and klx are clearly observed, along with the corresponding saturation points on each scale. A similar behavior at low illuminance can be observed for diffuser (b) in the right panel of Fig. 4, but only for values below klx, with notable prominence during sunset. For measurements within the range of values exceeding klx, the equipment does not undergo scale changes, resulting in continuous and seemingly anomaly-free measurements facilitated by diffuser (b). It is also evident from the graph that the ordinary white paper diffuser attenuates the signal to a greater extent compared to the tracing paper diffuser (as depicted on the right-hand side -axis of both plots). This difference can be attributed to the higher reflectivity of ordinary white paper, particularly within the visible region of the solar spectrum. Therefore, based on our findings, we recommend the use of diffuser (b) in this study. Custom selections may be done for other smartphones, however, this point requires special attention.
To validate this method, we compared it with reference equipment. In this case, the signal generated by the CHP1 pyrheliometer (in mV) is recorded by a Fisher Scientific DataTaker DT85 data logger and is converted to irradiance (in W/m2) through the equipment constant. This measurement is the reference DNI measurement of the LES lab, and it is recorded continuously on a minute scale as an average of instantaneous measurements every 15 s.
The precise measurement of DNI presents some difficulties. To carry it out, a pyrheliometer is used, an instrument that consists of an array of thermocouples (pyranometer) attached to a collimator tube, and a precision solar tracking mechanism. If the equipment is aligned with a precision of less than °, the pyrheliometer is capable of measuring the DNI with an uncertainty about 1. The measurements are carried out in broadband, that is, the irradiance corresponding to wavelengths between 200-4000 nm (which includes the entire solar spectrum) is integrated into a single value. Figure 2 shows the experimental setup of this work consisting of a Kipp & Zonen Zonen (2020) CHP1 pyrheliometer (blue oval) and a Samsung S5 smartphone (yellow oval) assembled in a Solys2 precision solar tracker. The assembly of the smartphone is shown in Fig. 3, in particular, its location perpendicular to the axis of the black bars and the assembly of the small hand-made collimator tube for the light sensor.
3.2 Calibration based on high precision clear-sky estimates
@@
As an alternative calibration method for places where a professional DNI measurement is not available, it is possible to use as reference accurate clear-sky estimates that use information from weather satellites and physically-consistent atmospheric models. This change in the reference implies a slight increase in the uncertainty in the determination of the calibration constants, since the DNI satellite estimate presents greater uncertainty than a ground reference measurement. There are sophisticated clear-sky models that integrate estimates of different atmospheric variables, either by satellite or by atmospheric reanalysis models, which can be considered as reference Gueymard (2008); Lefèvre et al. (2013) if they have been validated by terrestrial measurements with good concordance in various parts of the world.
@@
One interesting choice is the CAMS Programme (2020) (Copernicus Atmosphere Monitoring Service) platform which provides free clear-sky estimates using one of these reference models for the entire globe, the McClear model Lefèvre et al. (2013). This model is based on sophisticated radiative transfer calculations from the libRadTran Mayer and Kylling (2005) library and its operational version takes the form of a multiple input table based on real-time information on the state of the atmosphere. In particular, this model uses information on aerosols, precipitable water column, and ozone obtained from the CAMS reanalysis database and Earth albedo estimates obtained by the MODIS low-orbit satellite. The CAMS reanalysis in turn assimilates weather satellite’s information to provide its modeled data. Its platform enables access to 1-minute (and other time scales) solar irradiance estimates (global, direct and diffuse) from this high-precision model by simply entering the latitude and longitude of the site of interest. The file header contains information on each of the solar magnitudes provided. For example, the dimensions of radiation are Wh m-2 (irradiation, energy in the time interval per unit area), which must be converted to W m-2 (average power per unit area) by the corresponding conversion depending on the time scale.
4 Practical use of the smartphone light sensor
@@
will be used, where is the illuminance measured by the smartphone expressed in klx, are the DNI data expressed in W m-2, and and are two conversion constants to adjust. The calibration is performed with the two reference DNI data sets considered, the professional measurement of the pyrheliometer and the estimates of the McClear model. Minute measurements and estimates of November 24, 2021 were used at the LES experimental site, where clear-sky conditions were maintained throughout the day. Observing the plot of panel b in Fig. 4, for the adjustment we used only the data that meet klx, so that the smartphone sensor was always on the same scale of measurement. This value was chosen conservatively, in order to ensure measurements at intermediate values on the scale. The calibration constants for both cases are presented in Table 1 and the experimental fit in Fig. 5.
@@
It is interesting to compare the calibration curves. Satellite estimates exhibit deviations from the terrestrial measurements, so the calibration curve based on these data will present more uncertainty than the one obtained by comparison with ground measurements. As observed in Fig. 4, the McClear estimates for that day overestimate the direct irradiance. This leads the calibration curve obtained with this data set results above the calibration curve obtained with terrestrial measurements as observed in Fig. 5. Comparison of the McClear calibration curve with the measured DNI data reports a mean deviation of +2.1% (overestimation) and a mean square deviation of 2.6%. This uncertainty is above the measurement uncertainty of the reference instrument (1%), so it is distinguishable, but at the same time it is below the typical uncertainty of clear-sky satellite models (3-6%) Laguarda et al. (2020). This demonstrates that it is possible to perform the calibration based on satellite data of solar irradiance with a low uncertainty, enabling its use in the absence of reference measurements. Figure 6 shows the measurements obtained with the smartphone using both calibrations. As can be seen, in both cases a very good DNI measurement is achieved and completely acceptable for an instructional laboratory for a wide range of the day. The only downside is that the smartphone does not achieve a good measurement in the first and last minutes of the day where the illuminance is very low. Outside the diurnal range, i.e. when the sensor measurement is zero, the DNI measurement is affected by the non-zero offset of Eq. 8.
@@
These experimental results reveal that the smartphone is an outstanding tool for measuring solar irradiance through its illuminance sensor, applicable in low-cost university physics laboratories. Furthermore, the measurement capacity achieved with the equipment is really good, even in comparison to commercial sensors. Two relevant questions regarding the measurement capacity of the smartphone arises here. The first is to evaluate the typical uncertainty of a smartphone sensor used to measure DNI with both calibration methods. To answer this question would require data acquisition for several consecutive days (2-3 weeks) similar to professional calibrations following international standards ISO 9847:1992 (1992). The second is to evaluate the stability of the calibration curve over time, that is, how robust is the sensor to gradual degradation. This would require a professional calibration of the sensor every 3 months for about a year. With this set of tests it is possible to technically evaluate the capacity limits of the smartphone sensor for the solar irradiance measurement. In fact, moderate-cost professional irradiance sensors are recommended to be calibrated once a year, and a similar recommendation can arise for the smartphone sensor. We recommend here to perform the smartphone calibration each time before its utilization (i.e. the first day of measurements). Similar studies can be carried out for the measurement of global irradiance in the horizontal plane, not only direct irradiance, which surely requires evaluating the non-planar angular response of the smartphone sensor.